OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? a: (1, 7) b: (5, 0) c: (2, -1) d: (0, 5)

OpenStudy (anonymous):

Since it already tells you what y equals you just plug in equation one to equation 2 y = -3x + 5 and y = 2x + 5? \[\huge~\rm~-3x+5=2x+5\] Solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it's c?

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

please solve for x

OpenStudy (anonymous):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Moe95 Since it already tells you what y equals you just plug in equation one to equation 2 y = -3x + 5 and y = 2x + 5? \[\huge~\rm~-3x+5=2x+5\] Solve for x \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

d

OpenStudy (anonymous):

yep thats correct :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :)